OpenStudy (anonymous):

Find th an equation of a line that passes through (-4,-2) and has slope=0. Write the equation in both point-slope form and slope intercept form.

jhonyy9 (jhonyy9):

how do you think if there are x=-4 and y=-2 ?

OpenStudy (anonymous):

is it -4,0 and-2,0 with slope intercept y=4x-2

OpenStudy (anonymous):

(-4,-2)_m=0 Find the value of b using the formula for the equation of a line. y=mx+b Substitute the value of y into the equation. (-2)=(0)*(-4)+b Find the value of b. b=-2 Now that the values of m and b are known, substitute them into y=mx+b to find the equation of the line. y=-2

OpenStudy (anonymous):

thanks